# Cislunar Space Situational Awareness Constellation Design and Planning with Facility Location Problem

### v3

@@ 5.2 Lagrangian Heuristics Reallocation
1:$\bar{X}$, $\mathcal{J}^{\bar{Y}}$, $M^{b}$
2:$\underline{X}\leftarrow\bar{X}$triangleright Initialize allocation variable $X$ of feasible solution
3:$\underline{\theta}\leftarrow\boldsymbol{0}_{\ell\times q}$triangleright Initialize slack variable $\theta$ of feasible solution
@@
5:     $\mathcal{K}^{0}_{t}\leftarrow\mathcal{K}$triangleright Initialize set of unobserved targets at time $t$
6:     $\mathcal{J}^{0}_{t}\leftarrow\varnothing$triangleright Initialize set of observers with infeasible allocations
@@
9:              $\mathcal{K}^{0}_{t}\leftarrow\mathcal{K}^{0}_{t}\setminus k$triangleright Remove target $k$ from $\mathcal{K}^{0}_{t}$
@@
11:              $\mathcal{J}^{0}_{t}\leftarrow\mathcal{J}^{0}_{t}\cup\{j\}$triangleright Append violating facility index $j$ to $\mathcal{J}^{0}_{t}$
12:              $\underline{X}_{:jt}\leftarrow 0$triangleright Turn all allocations of facility $j$ at time $t$ off
@@
15:     $\underline{X}\leftarrow$ GreedyAllocation$(\underline{X},M^{b},\mathcal{J}^{0}_{t},\mathcal{K}^{0}_{t})$triangleright Perform allocations of facilities in $\mathcal{J}^{0}_{t}$
@@
18:              $\underline{\theta}_{tk}\leftarrow 1$triangleright If at least one observer is seeing target $k$ at time $t$, turn $\underline{\theta}_{tk}$ on
@@ 5.2.1 Greedy Allocation
4:     $\underline{X}_{i^{\star}j^{\star}t}\leftarrow 1$triangleright Turn allocation on
5:     $\mathcal{J}^{0}_{t}\leftarrow\mathcal{J}^{0}_{t}\setminus j^{\star}$triangleright Remove $j^{\star}$ from set of observers with infeasible allocations
6:     $\mathcal{K}^{0}_{t}\leftarrow\mathcal{K}^{0}_{t}\setminus\{k\mid M^{b}_{i^{\star}j^{\star}tk}=1\}$triangleright Remove newly targets observed from set of unobserved targets
@@ 5.2.2 Full-Factorial Allocation
1:$t,\underline{X}$, $M^{b}$, $\mathcal{J}^{0}_{t}$, $\mathcal{K}^{0}_{t}$
2:$\Pi_{\mathcal{J}}\leftarrow$permutations$(\mathcal{J}^{0}_{t})$triangleright Compute permutations of $j$ in $\mathcal{J}^{0}_{t}$
@@
4:     $\mathcal{K}^{0\pi}_{t}\leftarrow\mathcal{K}^{0}_{t}$triangleright Make a copy of $\mathcal{K}^{0}_{t}$ for current permutation
5:     $\underline{X}^{\pi}\leftarrow\underline{X}$triangleright Make a copy of $\underline{X}$ for current permutation
@@
14:                  $\mathcal{K}^{0\pi}_{t}\leftarrow\mathcal{K}^{0}_{t}\setminus\{k\mid M^{b}_{i^{\star}j^{\star}tk}=1\}$triangleright Remove newly targets observed from set of unobserved targets
@@
18:     $\Delta^{\pi}\leftarrow|\mathcal{K}^{0}_{t}|-|\mathcal{K}^{0\pi}_{t}|$triangleright Number of additional targets observed by current permutation
@@
20:$\pi^{*}\leftarrow\min_{\pi}\Delta^{\pi}$triangleright Choose allocation best permutation
21:$\underline{X}\leftarrow\underline{X}^{\pi^{*}}$triangleright Take $\underline{X}$ corresponding to best permutation
@@ 5.3 Heuristics for Neighborhood Facility Swap
1:$C_{j}$, $X_{\mathrm{best}}$, $Y_{\mathrm{best}}$, $Z_{\mathrm{best}}$, $\theta_{\mathrm{best}}$
2:$\mathcal{J}^{Y}\leftarrow\left\{j\in 1,\ldots,n\mid{Y_{\mathrm{best}}}_{j}=1\right\}$
@@
5:         $\mathcal{J}^{Y_{\mathrm{new}}}\leftarrow\mathcal{J}^{Y}\cup\{j_{\mathrm{in}}\}\setminus j_{\mathrm{out}}$triangleright Create new set of facilities by removing $j_{\mathrm{out}}$ and adding $j_{\mathrm{in}}$
6:         $Y_{\mathrm{new}}\leftarrow$ SetToVector$(\mathcal{J}^{Y_{\mathrm{new}}})$triangleright Convert set to vector
7:         $X_{\mathrm{new}},\theta_{\mathrm{new}}\leftarrow$ HeuristicFeasibleAllocation$(Y_{\mathrm{new}})$triangleright Heuristically compute feasible allocation
@@ 5.3.2 Inter-Orbit Neighborhood
The inter-orbit neighborhood of $j$ is defined as the set of candidate facility locations that are (1) on different LPOs with the same $M$:$N$ resonance, and (2) exhibit a similar solar phase angle $\phi_{S}$ with respect to a reference target at a reference epoch. The inter-orbit neighborhood for the same original slot as in Figure 9(a) is shown in Figure 9(b). The latter condition is required to isolate a single slot from the slots along an LPO with the same resonance as the LPO along which slot $j$ is located.
To arrive at an expression for the inter-orbit neighborhood of $j$, we begin by defining by $\mathcal{P}_{j}$ the set of LPOs that have the same $M$:$N$ resonance as the LPO of $j$. We also define a reference target position $\boldsymbol{r}_{k_{\rm ref}}$; in this work, we choose the mean location across all targets, $\boldsymbol{r}_{k_{\rm ref}}=\sum_{k}\boldsymbol{r}_{k}$. Then, for each candidate facility location $\xi$ along each LPO $\rho\in\mathcal{P}_{j}$, we determine the slot $\xi^{*}_{\rho}$ that achieves the closest solar phase angle with respect to $\boldsymbol{r}_{k_{\rm ref}}$ as the original slot $j$, given by
[…]
@@ 6 Numerical Results
Number of facility locations $n$ 1212 1212
@@
FOV $=60^{\circ}$ $\bar{m}_{\mathrm{crit}}=18$ 0.0460 0.0043
@@
FOV $=120^{\circ}$ $\bar{m}_{\mathrm{crit}}=18$ 0.1711 0.0176
@@ 6.1 Comparison of Lagrangian Method against Branch-and-Bound
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ diff. to $\Theta_{\mathrm{3600}}$
2 0.0419 101 OPTIMAL - - - 0.0402 620 $-3.99$ $-3.99$
3 0.0608 501 TIME_LIMIT 0.0608 681 OPTIMAL 0.0533 575 $-12.22$ $-12.22$
4 0.0778 500 TIME_LIMIT 0.0783 3601 TIME_LIMIT 0.0717 614 $-7.89$ $-8.54$
5 0.0939 500 TIME_LIMIT 0.0939 3601 TIME_LIMIT 0.0824 603 $-12.29$ $-12.29$
@@
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ diff. to $\Theta_{\mathrm{3600}}$
2 0.6058 501 TIME_LIMIT 0.7259 3601 TIME_LIMIT 0.6341 542 $+4.67$ $-12.65$
3 0.0000 501 TIME_LIMIT 0.6355 3601 TIME_LIMIT 0.7786 584 $+Inf$ $+22.52$
4 0.6756 501 TIME_LIMIT 0.6975 3601 TIME_LIMIT 0.8670 634 $+28.32$ $+24.31$
5 0.0000 500 TIME_LIMIT 0.7576 3602 TIME_LIMIT 0.9142 703 $+Inf$ $+20.68$
@@
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ diff. to $\Theta_{\mathrm{3600}}$
2 0.0000 501 TIME_LIMIT 0.6180 3602 TIME_LIMIT 0.6328 536 $+Inf$ $+2.39$
3 0.0000 501 TIME_LIMIT 0.8577 3601 TIME_LIMIT 0.7223 593 $+Inf$ $-15.79$
4 0.0000 501 TIME_LIMIT 0.7854 3602 TIME_LIMIT 0.9271 597 $+Inf$ $+18.04$
5 0.0000 501 TIME_LIMIT 0.9558 3603 TIME_LIMIT 0.9599 584 $+Inf$ $+0.43$
@@
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ diff. to $\Theta_{\mathrm{3600}}$
2 0.3737 66 OPTIMAL - - - 0.3144 523 $-15.86$ $-15.86$
3 0.3246 501 TIME_LIMIT 0.5078 1426 OPTIMAL 0.3967 644 $+22.20$ $-21.88$
4 0.1703 501 TIME_LIMIT 0.5973 1846 OPTIMAL 0.4681 519 $+174.92$ $-21.63$
5 0.1845 500 TIME_LIMIT 0.6670 3601 TIME_LIMIT 0.5471 537 $+196.50$ $-17.97$
@@
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ diff. to $\Theta_{\mathrm{3600}}$
2 0.9309 500 TIME_LIMIT 0.9398 2263 OPTIMAL 0.8339 638 $-10.42$ $-11.27$
3 0.0000 501 TIME_LIMIT 0.0000 3600 TIME_LIMIT 0.9056 634 $+Inf$ $+Inf$
4 0.0000 500 TIME_LIMIT 0.0000 3600 TIME_LIMIT 0.9355 524 $+Inf$ $+Inf$
5 0.0000 500 TIME_LIMIT 0.0000 3600 TIME_LIMIT 0.9461 569 $+Inf$ $+Inf$
@@
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ diff. to $\Theta_{\mathrm{3600}}$
2 0.4778 500 TIME_LIMIT 0.9701 3601 TIME_LIMIT 0.8681 526 $+81.70$ $-10.51$
3 0.4780 501 TIME_LIMIT 0.4780 3600 TIME_LIMIT 0.9461 537 $+97.91$ $+97.91$
4 0.4794 500 TIME_LIMIT 0.9957 819 OPTIMAL 0.9868 547 $+105.84$ $-0.89$
5 0.4794 500 TIME_LIMIT 0.9993 551 OPTIMAL 0.9774 560 $+103.88$ $-2.20$
@@ 6.2.1 Cone of Shame
The sensor steering history provides insights as to why in the $p=2$ case, the solution consists of one observer on the L1 Lyapunov orbit and another on a different LPO, while in cases with $p=3,4,5$ all observers placed on the L1 Lyapunov orbit. With at least $p=3$ observers, each can be tasked to observe at elevation angles of $0^{\circ}$, $-30^{\circ}$, and $30^{\circ}$ thereby complementing the observation to provide holistic coverage. Meanwhile, with only $p=2$ observers, the 3:1 resonant Halo, with higher cadence steering across azimuth and elevation angles, is found to be able to better complement the observation of a single L1 Lyapunov orbit observer. We note that the elevation angle spread for $p=3$ as seen in Figure 10(b) is not as thorough as with cases with $p=4$ and $5$, as seen in Figures 10(c) and 10(d). This is the reason why the LM solution is $-15.79\%$ worse than the B&B solution after 3600 seconds, as reported in Table 5(c), compared to cases with $p=2,4,5$.

### v4

@@ 5.2 Lagrangian Heuristics Reallocation
[…]
@@
5:  $\mathcal{K}^{0}_{t}\leftarrow\mathcal{K}$ $\triangleright$ Initialize set of unobserved targets at time $t$
6:  $\mathcal{J}^{0}_{t}\leftarrow\varnothing$ $\triangleright$ Initialize set of observers with infeasible allocations
@@
9:     $\mathcal{K}^{0}_{t}\leftarrow\mathcal{K}^{0}_{t}\setminus k$ $\triangleright$ Remove target $k$ from $\mathcal{K}^{0}_{t}$
@@
11:     $\mathcal{J}^{0}_{t}\leftarrow\mathcal{J}^{0}_{t}\cup\{j\}$ $\triangleright$ Append violating facility index $j$ to $\mathcal{J}^{0}_{t}$
12:     $\underline{X}_{:jt}\leftarrow 0$ $\triangleright$ Turn all allocations of facility $j$ at time $t$ off
@@
[…]
@@
18:     $\underline{\theta}_{tk}\leftarrow 1$ $\triangleright$ If at least one observer is seeing target $k$ at time $t$, turn $\underline{\theta}_{tk}$ on
@@ 5.2.1 Greedy Allocation
4:  $\underline{X}_{i^{\star}j^{\star}t}\leftarrow 1$ $\triangleright$ Turn allocation on
5:  $\mathcal{J}^{0}_{t}\leftarrow\mathcal{J}^{0}_{t}\setminus j^{\star}$ $\triangleright$ Remove $j^{\star}$ from set of observers with infeasible allocations
6:  $\mathcal{K}^{0}_{t}\leftarrow\mathcal{K}^{0}_{t}\setminus\{k\mid M^{b}_{i^{\star}j^{\star}tk}=1\}$ $\triangleright$ Remove newly targets observed from set of unobserved targets
@@ 5.2.2 Full-Factorial Allocation
[…]
@@
4:  $\mathcal{K}^{0\pi}_{t}\leftarrow\mathcal{K}^{0}_{t}$ $\triangleright$ Make a copy of $\mathcal{K}^{0}_{t}$ for current permutation
5:  $\underline{X}^{\pi}\leftarrow\underline{X}$ $\triangleright$ Make a copy of $\underline{X}$ for current permutation
@@
14:      $\mathcal{K}^{0\pi}_{t}\leftarrow\mathcal{K}^{0}_{t}\setminus\{k\mid M^{b}_{i^{\star}j^{\star}tk}=1\}$ $\triangleright$ Remove newly targets observed from set of unobserved targets
@@
18:  $\Delta^{\pi}\leftarrow|\mathcal{K}^{0}_{t}|-|\mathcal{K}^{0\pi}_{t}|$ $\triangleright$ Number of additional targets observed by current permutation
@@
[…]
@@ 5.3 Heuristics for Neighborhood Facility Swap
[…]
@@
5:   $\mathcal{J}^{Y_{\mathrm{new}}}\leftarrow\mathcal{J}^{Y}\cup\{j_{\mathrm{in}}\}\setminus j_{\mathrm{out}}$ $\triangleright$ Create new set of facilities by removing $j_{\mathrm{out}}$ and adding $j_{\mathrm{in}}$
[…]
@@ 5.3.2 Inter-Orbit Neighborhood
The inter-orbit neighborhood of $j$ is defined as the set of candidate facility locations that are (1) on different LPOs with the same $M$:NN resonance, and (2) exhibit a similar solar phase angle $\phi_{S}$ with respect to a reference target at a reference epoch. The inter-orbit neighborhood for the same original slot as in Figure 9(a) is shown in Figure 9(b). The latter condition is required to isolate a single slot from the slots along an LPO with the same resonance as the LPO along which slot $j$ is located.
To arrive at an expression for the inter-orbit neighborhood of $j$, we begin by defining by $\mathcal{P}_{j}$ the set of LPOs that have the same $M$:NN resonance as the LPO of $j$. We also define a reference target position $\boldsymbol{r}_{k_{\rm ref}}$; in this work, we choose the mean location across all targets, $\boldsymbol{r}_{k_{\rm ref}}=\sum_{k}\boldsymbol{r}_{k}$. Then, for each candidate facility location $\xi$ along each LPO $\rho\in\mathcal{P}_{j}$, we determine the slot $\xi^{*}_{\rho}$ that achieves the closest solar phase angle with respect to $\boldsymbol{r}_{k_{\rm ref}}$ as the original slot $j$, given by
$\xi^{*}_{\rho}=\underset{\xi}{\operatorname{argmin}}\left(\boldsymbol{l}_{jk_{\rm ref}}^{T}\boldsymbol{l}_{Sk_{\rm ref}}-\boldsymbol{l}_{\xi k_{\rm ref}}^{T}\boldsymbol{l}_{Sk_{\rm ref}}\right)$ (48)
@@ 6 Numerical Results
Number of facility locations $n$ $12$ $12$
@@
FOV $60^{\circ}$ $\bar{m}_{\mathrm{crit}}=18$ 0.0460 0.0043
@@
FOV $120^{\circ}$ $\bar{m}_{\mathrm{crit}}=18$ 0.1711 0.0176
@@ 6.1 Comparison of Lagrangian Method against Branch-and-Bound
$p$ B&B (TimeLimit =500=500) B&B (TimeLimit =3600=3600) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ % diff. to $\Theta_{\mathrm{3600}}$
2 0.0419 101 OPTIMAL - - - 0.0402 620 3.99-3.99 3.99-3.99
3 0.0608 501 TIME_LIMIT 0.0608 681 OPTIMAL 0.0533 575 12.22-12.22 12.22-12.22
4 0.0778 500 TIME_LIMIT 0.0783 3601 TIME_LIMIT 0.0717 614 7.89-7.89 8.54-8.54
5 0.0939 500 TIME_LIMIT 0.0939 3601 TIME_LIMIT 0.0824 603 12.29-12.29 12.29-12.29
@@
$p$ B&B (TimeLimit =500=500) B&B (TimeLimit =3600=3600) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ % diff. to $\Theta_{\mathrm{3600}}$
2 0.6058 501 TIME_LIMIT 0.7259 3601 TIME_LIMIT 0.6341 542 +4.67+4.67 12.65-12.65
3 0.0000 501 TIME_LIMIT 0.6355 3601 TIME_LIMIT 0.7786 584 +Inf+Inf +22.52+22.52
4 0.6756 501 TIME_LIMIT 0.6975 3601 TIME_LIMIT 0.8670 634 +28.32+28.32 +24.31+24.31
5 0.0000 500 TIME_LIMIT 0.7576 3602 TIME_LIMIT 0.9142 703 +Inf+Inf +20.68+20.68
@@
$p$ B&B (TimeLimit =500=500) B&B (TimeLimit =3600=3600) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ % diff. to $\Theta_{\mathrm{3600}}$
2 0.0000 501 TIME_LIMIT 0.6180 3602 TIME_LIMIT 0.6328 536 +Inf+Inf +2.39+2.39
3 0.0000 501 TIME_LIMIT 0.8577 3601 TIME_LIMIT 0.7223 593 +Inf+Inf 15.79-15.79
4 0.0000 501 TIME_LIMIT 0.7854 3602 TIME_LIMIT 0.9271 597 +Inf+Inf +18.04+18.04
5 0.0000 501 TIME_LIMIT 0.9558 3603 TIME_LIMIT 0.9599 584 +Inf+Inf +0.43+0.43
@@
$p$ B&B (TimeLimit =500=500) B&B (TimeLimit =3600=3600) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ % diff. to $\Theta_{\mathrm{3600}}$
2 0.3737 66 OPTIMAL - - - 0.3144 523 15.86-15.86 15.86-15.86
3 0.3246 501 TIME_LIMIT 0.5078 1426 OPTIMAL 0.3967 644 +22.20+22.20 21.88-21.88
4 0.1703 501 TIME_LIMIT 0.5973 1846 OPTIMAL 0.4681 519 +174.92+174.92 21.63-21.63
5 0.1845 500 TIME_LIMIT 0.6670 3601 TIME_LIMIT 0.5471 537 +196.50+196.50 17.97-17.97
@@
$p$ B&B (TimeLimit =500=500) B&B (TimeLimit =3600=3600) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ % diff. to $\Theta_{\mathrm{3600}}$
2 0.9309 500 TIME_LIMIT 0.9398 2263 OPTIMAL 0.8339 638 10.42-10.42 11.27-11.27
3 0.0000 501 TIME_LIMIT 0.0000 3600 TIME_LIMIT 0.9056 634 +Inf+Inf +Inf+Inf
4 0.0000 500 TIME_LIMIT 0.0000 3600 TIME_LIMIT 0.9355 524 +Inf+Inf +Inf+Inf
5 0.0000 500 TIME_LIMIT 0.0000 3600 TIME_LIMIT 0.9461 569 +Inf+Inf +Inf+Inf
@@
$p$ B&B (TimeLimit =500=500) B&B (TimeLimit =3600=3600) Lagrangian Method
$\Theta_{\mathrm{500}}$ Time Status $\Theta_{\mathrm{3600}}$ Time Status $\Theta_{\mathrm{LH}}$ Time % diff. to $\Theta_{\mathrm{500}}$ % diff. to $\Theta_{\mathrm{3600}}$
2 0.4778 500 TIME_LIMIT 0.9701 3601 TIME_LIMIT 0.8681 526 +81.70+81.70 10.51-10.51
3 0.4780 501 TIME_LIMIT 0.4780 3600 TIME_LIMIT 0.9461 537 +97.91+97.91 +97.91+97.91
4 0.4794 500 TIME_LIMIT 0.9957 819 OPTIMAL 0.9868 547 +105.84+105.84 0.89-0.89
5 0.4794 500 TIME_LIMIT 0.9993 551 OPTIMAL 0.9774 560 +103.88+103.88 2.20-2.20
@@ 6.2.1 Cone of Shame
The sensor steering history provides insights as to why in the $p=2$ case, the solution consists of one observer on the L1 Lyapunov orbit and another on a different LPO, while in cases with $p=3,4,5$ all observers placed on the L1 Lyapunov orbit. With at least $p=3$ observers, each can be tasked to observe at elevation angles of $0^{\circ}$, $30^{\circ}$, and $30^{\circ}$ - thereby complementing the observation to provide holistic coverage. Meanwhile, with only $p=2$ observers, the 3:1 resonant Halo, with higher cadence steering across azimuth and elevation angles, is found to be able to better complement the observation of a single L1 Lyapunov orbit observer. We note that the elevation angle spread for $p=3$ as seen in Figure 10(b) is not as thorough as with cases with $p=4$ and $5$, as seen in Figures 10(c) and 10(d). This is the reason why the LM solution is $15.79\%$ worse than the B&B solution after 3600 seconds, as reported in Table 5(c), compared to cases with $p=2,4,5$.